# Agentic Auto-Scheduling: An Experimental Study of LLM-Guided Loop Optimization

### v1

@@ I Introduction
Loop nests are critical performance bottlenecks, particularly in scientific computing, image processing, and machine learning domains. Decades of research have yielded powerful optimization techniques, from compiler heuristics (e.g., in GCC [1] and LLVM [2]) and sophisticated polyhedral methods [3] (e.g., Pluto [4]) that analyze dependencies and orchestrate complex transformations (like tiling, fusion, parallelization), to empirical autotuning frameworks [5, 6, 7]. Despite these advances, consistently achieving optimal performance across diverse applications and hardware remains challenging, motivating the exploration of complementary optimization strategies.
@@
Recent advancements in Large Language Models (LLMs) open up intriguing new possibilities. Beyond their well-known text generation capabilities, LLMs exhibit remarkable abilities in understanding complex instructions, reasoning about problems, and even analyzing source code structure. This potential has spurred recent investigations into using LLMs for code optimization tasks [8, 9, 10, 11, 12, 13, 14]. However, approaches generating optimized code directly [12, 13, 11, 14] often struggle to guarantee the semantic correctness of complex transformations without costly formal verification or brittle output checking. Alternatively, other works focus on selecting compiler passes or flags [8, 9, 10]. These methods are useful for orchestrating existing compiler heuristics, often at the Intermediate Representation (IR) level. However, they typically lack fine-grained control for composing sequences of high-level source transformations which are often essential for maximizing data locality and parallelism gains. Additionally, prior works often prioritized objectives other than execution speed, such as code size, and necessitated domain-specific LLM fine-tuning. These challenges motivate exploring a more agentic paradigm, where the LLM acts as a proactive decision-maker interacting with its environment. This leads to our central research question: Can off-the-shelf LLMs, grounded by empirical compiler feedback, effectively guide the complex process of loop optimization?
@@
To explore this question, we introduce ComPilot (Compiler Pilot), an experimental framework built around a closed-loop interaction between an LLM and a compiler infrastructure. In ComPilot, the LLM acts as an optimization agent, iteratively proposing sequences of loop transformations (schedules) for a given loop nest. These schedules are then passed to a compiler infrastructure (in our case, the Tiramisu compiler [15]) which attempts to apply the transformations, checks their legality based on dependence analysis [16, 17], and generates code. ComPilot incorporates a feedback loop: the outcome of applying a schedule (success, failure type, and measured speedup/slowdown if successful) is reported back to the LLM. This closed-loop interaction allows the LLM to potentially learn from its successes and failures and refine its optimization strategy based on direct empirical evidence from the target machine. This approach leverages LLM capabilities to explore the transformation space while relying on the compiler’s rigor for legality checking and code generation, all without requiring LLM fine-tuning or direct generation of transformed code.
@@
A demonstration of its effectiveness on PolyBench [18], showing that off-the-shelf LLMs can achieve significant speedups over strong baselines by discovering high-performance sequences of transformations.
@@
Our evaluation demonstrates the potential of LLM-guided optimization. Using a best-of-5 runs strategy, our approach achieves a geometric mean speedup of 3.54 over the original code and 2.94 over the state-of-the-art Pluto polyhedral compiler across the PolyBench suite [18]. On certain benchmarks, ComPilot discovers schedules yielding speedups exceeding 100 compared to the original code.
@@ II-A Context Initialization Phase
Finally, the Context Initializer Module prompts the LLM to first analyze the target loop nest before proposing transformations. An example of such an analysis is shown in Figure 4. This analysis serves as a form of chain-of-thought reasoning [19], guiding all the LLM’s subsequent optimization decisions. The LLM is expected to: deconstruct the loop nest’s structure, infer the purpose of the loop nest and the role of individual computations, and propose potential optimization strategies based on its understanding. The importance of this program analysis phase is demonstrated experimentally in the appendix (RQ10).
@@ II-B Iterative Optimization Phase
Schedules that pass this initial check proceed to a formal legality check performed by the backend compiler. In our implementation, Tiramisu employs rigorous polyhedral dependence analysis [16, 17] to guarantee that a transformation preserves the original program’s semantics. This delegation is a core design principle: ComPilot leverages the LLM for high-level strategic exploration while entrusting the compiler with formal correctness, ensuring code reliability without brittle runtime output comparisons. For legal schedules, the Response Parser emits the corresponding Tiramisu API calls to implement the transformations. When parameter calculations are required (e.g., skewing factors or fusion-induced shifting), ComPilot invokes Tiramisu’s internal solvers. The transformed code is then compiled and executed on the target machine to measure runtime and compute speedup or slowdown.
@@
This feedback, containing the outcome (success/failure type, performance metrics), is appended to the optimization history to update the LLM’s working memory. For the next iteration, the LLM processes this entire updated context. This makes the LLM utilize its inherent in-context learning (ICL) [20] capabilities to interpret this feedback and the preceding interaction history, adapting its strategy and reasoning process to generate the subsequent schedule proposition. This mechanism allows ComPilot to leverage the LLM’s adaptability dynamically within the dialogue, without requiring explicit fine-tuning or gradient updates, relying solely on the information provided in the prompt context.
@@ III Results and Analysis
Benchmarks: We evaluated ComPilot on the PolyBench/C benchmark suite 111https://polybench.sf.net [18](version 4.2.1), a standard benchmark set for polyhedral compilation research. PolyBench comprises 30 benchmarks from diverse domains (linear algebra, stencils, etc.). For each benchmark, we used all five standard dataset sizes and default data types, resulting in 150 distinct benchmark instances.
@@ III-A Main Results
To better contextualize the performance gains reported in RQ1, we compare ComPilot against a recognized state-of-the-art polyhedral optimizer, Pluto [4]. Pluto employs a heuristic-driven approach to optimize for parallelism and data locality based on analyzing data dependencies within loop nests. Unlike ComPilot, Pluto does not rely on execution feedback during its decision process.
@@
We also compare ComPilot against the deep-learning-based Tiramisu autoscheduler [21], restricting evaluation to the 8 of 30 PolyBench benchmarks it supports. As noted in subsequent work [22], this autoscheduler cannot handle non-rectangular iteration domains or programs with multiple loop nests—limitations that exclude most of PolyBench. On these 8 benchmarks, ComPilot shows a clear advantage: ComPilot@ attains a geometric mean speedup of 2.65 (95% CI: [2.52, 2.69]) over the autoscheduler, rising to 3.23 (95% CI: [3.14, 3.32]) with ComPilot@. Figure 12 reports per-benchmark and per-size details.
This gap stems from three design differences: (1) ComPilot explores a richer transformation space, including skewing and reversal (critical for stencil dependencies in jacobi2d and seidel2d) that the autoscheduler does not support; (2) it employs a more flexible exploration strategy, iteratively combining and refining schedules rather than applying a fixed order, non-repetitive sequence; and (3) it is guided by ground-truth empirical feedback, while [21] relies on predictions from an offline-trained cost model; this lets ComPilot adapt to the specific performance characteristics of the target hardware and avoid being misled by inaccurate model predictions.
@@ III-B Ablation Study
ComPilot’s iterative process relies heavily on providing feedback (legality, validity, measured speedup/slowdown) to the LLM after each transformation attempt. This feedback loop serves a purpose analogous to Retrieval-Augmented Generation (RAG) [23], providing external, factual context to ground the LLM’s next suggestion. However, unlike typical RAG where context comes from a static corpus, ComPilot’s context is dynamically generated from interaction with the compiler and execution environment. Here, we assess the importance of this dynamic feedback mechanism.
@@
ComPilot’s design intentionally separates concerns: the LLM provides high-level transformation guidance, while the compiler handles the complex tasks of applying transformations and ensuring legality via dependence analysis [16, 17]. To assess the benefits of this delegation, we evaluated an alternative approach where the LLM directly generates the transformed C code. We implemented a variant of the ComPilot workflow for this comparison. Instead of receiving transformation commands, this variant prompted the LLM to directly rewrite the input C kernel code extracted from PolyBench benchmarks. The goal was for the LLM to output optimized C code. The iterative dialogue structure remained, but the LLM’s output and the subsequent verification steps were fundamentally different.
@@ IV Related Work
LLMs have shown significant potential in software engineering tasks in recent years. Whether for code generation, code translation between programming languages, test case generation, or code summarization [24]. This has spurred significant interest in leveraging LLMs for compiler-related tasks, particularly in code transformation and optimization.
@@ IV-A Prior Work on LLM-Based Code Optimization
The Meta LLM Compiler [8] extends Code Llama [25] with additional pretraining on LLVM-IR and assembly to specialize in low-level tasks (e.g., predicting optimization pass outcomes), followed by extensive fine-tuning for downstream compiler tasks such as flag selection for LLVM IR code-size reduction; it also introduces PassListEval to validate pass lists via unit testing and detect semantic breaks or crashes. Cummins et al. [9] train a 7B-parameter LLVM-IR model to predict pass sequences for code-size reduction, relying on the compiler to apply transformations and ensure correctness. Grubisic et al. [10] add an iterative feedback loop that reports pass sequence validity and compilation errors, enabling the LLM to repair its outputs.
@@ IV-A2 High-Level Code Optimization
Rosas et al. [11] evaluate LLMs for OpenMP parallelization via direct code generation, validating correctness by checkpointing and comparing variable states against the original program. Shypula et al. [12] fine-tune on the Performance-Improving Edits (PIE) dataset curated from CodeNet [26], assessing correctness with unit tests, while Duan et al. [13] fine-tune CodeT5 [27] on PIE using reinforcement learning. LLM-Vectorizer [14] combines off-the-shelf GPT-4 with formal verification (Alive2 [28]) and a multi-agent finite-state machine to generate, validate, and repair vectorized code. Concurrently, CompilerGPT [29] employs an iterative LLM–compiler feedback loop in which the LLM analyzes Clang/GCC optimization reports and rewrites C++ to better trigger compiler heuristics, verifying correctness via a user-provided test harness.
@@ IV-B Positioning of Our Work
Prior LLM-based code optimization largely uses domain-specialized models trained from scratch or fine-tuned on compiler-centric data (e.g., LLVM-IR, assembly) or on performance-improving edits, typically to directly generate optimized code. When framed as flag selection, the objective is usually code-size reduction rather than execution speed. Correctness is commonly checked via output comparison (which cannot guarantee semantic preservation) or via formal verification, whose support for complex optimizations and scalability to large programs is limited. In contrast, we leverage general-purpose, off-the-shelf LLMs (no fine-tuning) to iteratively select and guide optimizations. A key distinction from CompilerGPT [29]—which analyzes compiler reports and rewrites source to better trigger compiler heuristics—is that our LLM emits compiler API calls for loop transformations, while the compiler applies them and ensures legality via dependence analysis. This offloads code generation and provides correctness guarantees, unlike direct code-generation models and CompilerGPT, which rely on unit tests (whose potential for correctness failures we demonstrate in RQ7). Table II summarizes these differences.
@@
| Feature | Ours | [8] | [9] | [10] | [11] | [12] | [13] | [14] |
@@ IV-C1 Polyhedral Compilers and Autoschedulers
The polyhedral model [3] is a mathematical model for representing code and code transformations and is used in state-of-the-art compilers to apply complex code transformations and reason about their correctness [30, 31, 32, 33, 34, 35, 15, 36, 37, 38, 39, 40, 41, 42, 43]. Tools such as Pluto [4], leverage the polyhedral model to represent loop nests enabling systematic exploration of loop transformations like tiling, fusion, and skewing. Pluto employs an integer linear programming (ILP) solver to derive affine transformations that optimize data locality and parallelism, guided by a linear objective function. PolyGym [44] extends this paradigm with a reinforcement learning environment to explore schedules.
@@ IV-C2 Non-Polyhedral Autoschedulers
Halide [45] and TVM [46] adopt domain-specific auto-scheduling with deep-learning cost models. Halide combines beam search with a feedforward neural network to predict execution times from handcrafted features (e.g., loop nesting depth, memory access patterns). Similarly, AutoTVM [47] uses gradient-based optimization to tune tensor operations for accelerators. These frameworks excel in domain-specific contexts (e.g., image processing, DNN kernels).

### v2

@@ I Introduction
Loop nests are critical performance bottlenecks, particularly in scientific computing, image processing, and machine learning domains. Decades of research have yielded powerful optimization techniques, from compiler heuristics (e.g., in GCC [stallman2002gnu] and LLVM [lattner2004llvm]) and sophisticated polyhedral methods [Feautrier2011] (e.g., Pluto [bondhugula2008practical]) that analyze dependencies and orchestrate complex transformations (like tiling, fusion, parallelization), to empirical autotuning frameworks [ansel2014opentuner, 8423171, Vuduc2011]. Despite these advances, consistently achieving optimal performance across diverse applications and hardware remains challenging, motivating the exploration of complementary optimization strategies.
@@
Recent advancements in Large Language Models (LLMs) open up intriguing new possibilities. Beyond their well-known text generation capabilities, LLMs exhibit remarkable abilities in understanding complex instructions, reasoning about problems, and even analyzing source code structure. This potential has spurred recent investigations into using LLMs for code optimization tasks [cummins2024meta, cummins2023large, grubisic2024compiler, rosas2024should, shypula2023learning, PerfRL, taneja2025llm]. However, approaches generating optimized code directly [shypula2023learning, PerfRL, rosas2024should, taneja2025llm] often struggle to guarantee the semantic correctness of complex transformations without costly formal verification or brittle output checking. Alternatively, other works focus on selecting compiler passes or flags [cummins2024meta, cummins2023large, grubisic2024compiler]. These methods are useful for orchestrating existing compiler heuristics, often at the Intermediate Representation (IR) level. However, they typically lack fine-grained control for composing sequences of high-level source transformations which are often essential for maximizing data locality and parallelism gains. Additionally, prior works often prioritized objectives other than execution speed, such as code size, and necessitated domain-specific LLM fine-tuning. These challenges motivate exploring a more agentic paradigm, where the LLM acts as a proactive decision-maker interacting with its environment. This leads to our central research question: Can off-the-shelf LLMs, grounded by empirical compiler feedback, effectively guide the complex process of loop optimization?
@@
To explore this question, we introduce ComPilot (Compiler Pilot), an experimental framework built around a closed-loop interaction between an LLM and a compiler infrastructure. In ComPilot, the LLM acts as an optimization agent, iteratively proposing sequences of loop transformations (schedules) for a given loop nest. These schedules are then passed to a compiler infrastructure (in our case, the Tiramisu compiler [tiramisu]) which attempts to apply the transformations, checks their legality based on dependence analysis [feautrier_array_1988, violateddep], and generates code. ComPilot incorporates a feedback loop: the outcome of applying a schedule (success, failure type, and measured speedup/slowdown if successful) is reported back to the LLM. This closed-loop interaction allows the LLM to potentially learn from its successes and failures and refine its optimization strategy based on direct empirical evidence from the target machine. This approach leverages LLM capabilities to explore the transformation space while relying on the compiler’s rigor for legality checking and code generation, all without requiring LLM fine-tuning or direct generation of transformed code.
@@
A demonstration of its effectiveness on PolyBench [louis-noel_polybench_2010], showing that off-the-shelf LLMs can achieve significant speedups over strong baselines by discovering high-performance sequences of transformations.
@@
Our evaluation demonstrates the potential of LLM-guided optimization. Using a best-of-5 runs strategy, our approach achieves a geometric mean speedup of 3.54 over the original code and 2.94 over the state-of-the-art Pluto polyhedral compiler across the PolyBench suite [louis-noel_polybench_2010]. On certain benchmarks, ComPilot discovers schedules yielding speedups exceeding 100 compared to the original code.
@@ II-A Context Initialization Phase
Finally, the Context Initializer Module prompts the LLM to first analyze the target loop nest before proposing transformations. An example of such an analysis is shown in Figure 4. This analysis serves as a form of chain-of-thought reasoning [NEURIPS2022_9d560961], guiding all the LLM’s subsequent optimization decisions. The LLM is expected to: deconstruct the loop nest’s structure, infer the purpose of the loop nest and the role of individual computations, and propose potential optimization strategies based on its understanding. The importance of this program analysis phase is demonstrated experimentally in the appendix (RQ10).
@@ II-B Iterative Optimization Phase
Schedules that pass this initial check proceed to a formal legality check performed by the backend compiler. In our implementation, Tiramisu employs rigorous polyhedral dependence analysis [feautrier_array_1988, violateddep] to guarantee that a transformation preserves the original program’s semantics. This delegation is a core design principle: ComPilot leverages the LLM for high-level strategic exploration while entrusting the compiler with formal correctness, ensuring code reliability without brittle runtime output comparisons. For legal schedules, the Response Parser emits the corresponding Tiramisu API calls to implement the transformations. When parameter calculations are required (e.g., skewing factors or fusion-induced shifting), ComPilot invokes Tiramisu’s internal solvers. The transformed code is then compiled and executed on the target machine to measure runtime and compute speedup or slowdown.
@@
This feedback, containing the outcome (success/failure type, performance metrics), is appended to the optimization history to update the LLM’s working memory. For the next iteration, the LLM processes this entire updated context. This makes the LLM utilize its inherent in-context learning (ICL) [NEURIPS2020_1457c0d6] capabilities to interpret this feedback and the preceding interaction history, adapting its strategy and reasoning process to generate the subsequent schedule proposition. This mechanism allows ComPilot to leverage the LLM’s adaptability dynamically within the dialogue, without requiring explicit fine-tuning or gradient updates, relying solely on the information provided in the prompt context.
@@ III Results and Analysis
Compiler: We utilized the Tiramisu compiler [tiramisu] (commit 041afad) to perform legality checks, invoke its internal solvers (e.g., for skewing factors), apply the LLM-suggested transformation sequences, and generate executable code.
@@
Benchmarks: We evaluated ComPilot on the PolyBench/C benchmark suite 111https://polybench.sf.net [louis-noel_polybench_2010](version 4.2.1), a standard benchmark set for polyhedral compilation research. PolyBench comprises 30 benchmarks from diverse domains (linear algebra, stencils, etc.). For each benchmark, we used all five standard dataset sizes and default data types, resulting in 150 distinct benchmark instances.
@@ III-A Main Results
To better contextualize the performance gains reported in RQ1, we compare ComPilot against a recognized state-of-the-art polyhedral optimizer, Pluto [bondhugula2008practical]. Pluto employs a heuristic-driven approach to optimize for parallelism and data locality based on analyzing data dependencies within loop nests. Unlike ComPilot, Pluto does not rely on execution feedback during its decision process.
@@
We also compare ComPilot against the deep-learning-based Tiramisu autoscheduler [baghdadi2021deep], restricting evaluation to the 8 of 30 PolyBench benchmarks it supports. As noted in subsequent work [looper], this autoscheduler cannot handle non-rectangular iteration domains or programs with multiple loop nests—limitations that exclude most of PolyBench. On these 8 benchmarks, ComPilot shows a clear advantage: ComPilot@ attains a geometric mean speedup of 2.65 (95% CI: [2.52, 2.69]) over the autoscheduler, rising to 3.23 (95% CI: [3.14, 3.32]) with ComPilot@. Figure 12 reports per-benchmark and per-size details.
This gap stems from three design differences: (1) ComPilot explores a richer transformation space, including skewing and reversal (critical for stencil dependencies in jacobi2d and seidel2d) that the autoscheduler does not support; (2) it employs a more flexible exploration strategy, iteratively combining and refining schedules rather than applying a fixed order, non-repetitive sequence; and (3) it is guided by ground-truth empirical feedback, while [baghdadi2021deep] relies on predictions from an offline-trained cost model; this lets ComPilot adapt to the specific performance characteristics of the target hardware and avoid being misled by inaccurate model predictions.
@@ III-B Ablation Study
ComPilot’s iterative process relies heavily on providing feedback (legality, validity, measured speedup/slowdown) to the LLM after each transformation attempt. This feedback loop serves a purpose analogous to Retrieval-Augmented Generation (RAG) [NEURIPS2020_6b493230], providing external, factual context to ground the LLM’s next suggestion. However, unlike typical RAG where context comes from a static corpus, ComPilot’s context is dynamically generated from interaction with the compiler and execution environment. Here, we assess the importance of this dynamic feedback mechanism.
@@
ComPilot’s design intentionally separates concerns: the LLM provides high-level transformation guidance, while the compiler handles the complex tasks of applying transformations and ensuring legality via dependence analysis [feautrier_array_1988, violateddep]. To assess the benefits of this delegation, we evaluated an alternative approach where the LLM directly generates the transformed C code. We implemented a variant of the ComPilot workflow for this comparison. Instead of receiving transformation commands, this variant prompted the LLM to directly rewrite the input C kernel code extracted from PolyBench benchmarks. The goal was for the LLM to output optimized C code. The iterative dialogue structure remained, but the LLM’s output and the subsequent verification steps were fundamentally different.
@@ IV Related Work
LLMs have shown significant potential in software engineering tasks in recent years. Whether for code generation, code translation between programming languages, test case generation, or code summarization [zheng2023survey]. This has spurred significant interest in leveraging LLMs for compiler-related tasks, particularly in code transformation and optimization.
@@ IV-A Prior Work on LLM-Based Code Optimization
The Meta LLM Compiler [cummins2024meta] extends Code Llama [roziere2023code] with additional pretraining on LLVM-IR and assembly to specialize in low-level tasks (e.g., predicting optimization pass outcomes), followed by extensive fine-tuning for downstream compiler tasks such as flag selection for LLVM IR code-size reduction; it also introduces PassListEval to validate pass lists via unit testing and detect semantic breaks or crashes. Cummins et al. [cummins2023large] train a 7B-parameter LLVM-IR model to predict pass sequences for code-size reduction, relying on the compiler to apply transformations and ensure correctness. Grubisic et al. [grubisic2024compiler] add an iterative feedback loop that reports pass sequence validity and compilation errors, enabling the LLM to repair its outputs.
@@ IV-A2 High-Level Code Optimization
Rosas et al. [rosas2024should] evaluate LLMs for OpenMP parallelization via direct code generation, validating correctness by checkpointing and comparing variable states against the original program. Shypula et al. [shypula2023learning] fine-tune on the Performance-Improving Edits (PIE) dataset curated from CodeNet [puri2021codenet], assessing correctness with unit tests, while Duan et al. [PerfRL] fine-tune CodeT5 [wang2021codet5] on PIE using reinforcement learning. LLM-Vectorizer [taneja2025llm] combines off-the-shelf GPT-4 with formal verification (Alive2 [lopes2021alive2]) and a multi-agent finite-state machine to generate, validate, and repair vectorized code. Concurrently, CompilerGPT [compilergpt] employs an iterative LLM–compiler feedback loop in which the LLM analyzes Clang/GCC optimization reports and rewrites C++ to better trigger compiler heuristics, verifying correctness via a user-provided test harness.
@@ IV-B Positioning of Our Work
Prior LLM-based code optimization largely uses domain-specialized models trained from scratch or fine-tuned on compiler-centric data (e.g., LLVM-IR, assembly) or on performance-improving edits, typically to directly generate optimized code. When framed as flag selection, the objective is usually code-size reduction rather than execution speed. Correctness is commonly checked via output comparison (which cannot guarantee semantic preservation) or via formal verification, whose support for complex optimizations and scalability to large programs is limited. In contrast, we leverage general-purpose, off-the-shelf LLMs (no fine-tuning) to iteratively select and guide optimizations. A key distinction from CompilerGPT [compilergpt]—which analyzes compiler reports and rewrites source to better trigger compiler heuristics—is that our LLM emits compiler API calls for loop transformations, while the compiler applies them and ensures legality via dependence analysis. This offloads code generation and provides correctness guarantees, unlike direct code-generation models and CompilerGPT, which rely on unit tests (whose potential for correctness failures we demonstrate in RQ7). Table II summarizes these differences.
@@
| Feature | Ours | [cummins2024meta] | [cummins2023large] | [grubisic2024compiler] | [rosas2024should] | [shypula2023learning] | [PerfRL] | [taneja2025llm] |
@@ IV-C1 Polyhedral Compilers and Autoschedulers
The polyhedral model [Feautrier2011] is a mathematical model for representing code and code transformations and is used in state-of-the-art compilers to apply complex code transformations and reason about their correctness [wolf1991loop, lefebvre_automatic_1998, Qui00, thies_unified_2001, Darte_contraction_2005, baghdadi2015PhD, tiramisu, trifunovic_graphite_2010, polly, tobias_hexagonal_cgo13, Vasilache2018TensorCF, baghdadi2011speculation, merouani2020deep, pouchet.11.popl, baghdadi2020tiramisuDNNDenseSparse]. Tools such as Pluto [bondhugula2008practical], leverage the polyhedral model to represent loop nests enabling systematic exploration of loop transformations like tiling, fusion, and skewing. Pluto employs an integer linear programming (ILP) solver to derive affine transformations that optimize data locality and parallelism, guided by a linear objective function. PolyGym [brauckmann2021polygym] extends this paradigm with a reinforcement learning environment to explore schedules.
@@ IV-C2 Non-Polyhedral Autoschedulers
Halide [adams2019learning] and TVM [chen2018tvm] adopt domain-specific auto-scheduling with deep-learning cost models. Halide combines beam search with a feedforward neural network to predict execution times from handcrafted features (e.g., loop nesting depth, memory access patterns). Similarly, AutoTVM [chen2018learning] uses gradient-based optimization to tune tensor operations for accelerators. These frameworks excel in domain-specific contexts (e.g., image processing, DNN kernels).